Please show me how to evaluate each expression:
a. 8^2/3
b. (-1)^1/5
c. (-27)^2/3

Thank you for your help

Answer by lwsshak3(11628)

Please show me how to evaluate each expression:
a. 8^2/3
Take the cube root of 8=3, then square it=3^2=9
..
b. (-1)^1/5
Take the 5th root of -1=-1. (note: the odd root of a negative number is also negative)
..
c. (-27)^2/3
Take the cube root of -27=-3, then square it (-3)^2=9